# Reliable Grid Forecasting: State Space Models for Safety-Critical Energy Systems

### v5

@@ Abstract
Using this framework, we evaluate state space models (Mamba variants) and strong baselines on a weather-aligned California Independent System Operator (CAISO) dataset spanning Nov 2023–Nov 2025 (84,498 hourly records across 5 regional transmission areas) under a rolling-origin walk-forward backtest. We develop and evaluate thermal-lag-aligned weather fusion strategies for these architectures.
Our results demonstrate that standard accuracy metrics are insufficient proxies for operational safety: models with comparable MAPE can imply materially different tail reserve requirements (Reserve). We show that explicit weather integration narrows error distributions, reducing the impact of temperature-driven demand spikes. Furthermore, while probabilistic calibration reduces large-error events, it can induce systematic schedule inflation. We introduce Bias/OPR-constrained objectives to enable auditable trade-offs between minimizing tail risk and preventing trivial over-forecasting.
keywords:
Load forecasting, state space models, Mamba, deep learning, probabilistic forecasting, California grid, CAISO1 Introduction
@@
We present a systematic evaluation of Mamba architectures for reliable California grid load forecasting. Benchmarking against iTransformer and the foundation model Chronos, we shift the evaluation focus from pure accuracy to operational safety. Our walk-forward results highlight that accuracy alone is not a sufficient proxy for operational risk: Reserve can vary substantially across models even when MAPE is comparable. Furthermore, we demonstrate that explicit weather integration is critical for safety by narrowing error distributions and reducing extreme error events (Section 6.3).
@@
Grid-Specific Evaluation Framework. We formalize operational risk metrics (Under-Prediction Rate, tail Reserve requirements, Bias/OPR diagnostics) that capture asymmetric operational costs and prevent “fake safety” via systematic forecast inflation (Section 2.3).
@@
Loss vs. Safety Analysis. We show that proper probabilistic calibration (multi-quantile pinball) can reduce large-error events on fixed evaluation splits, but can also introduce systematic inflation unless Bias/OPR are explicitly constrained and jointly reported with tail-risk metrics (Section 6).
2 Background
2.1 State Space Models
@@
2.2 Load Forecasting Fundamentals
@@
2.3 Grid-Specific Evaluation Metrics
@@
2.3.1 Operational Cost Asymmetry
@@
2.3.2 Proposed Metrics
- Under-Prediction Rate (UPR).
@@ 2.3.2 Proposed Metrics
The frequency of under-estimation events represents the probability of needing real-time upward dispatch:
@@
(5) A naive model might achieve low MAPE by constantly under-predicting (taking the median); UPR exposes this risky behavior.
- Over-Prediction Rate (OPR).
@@
The frequency of over-forecasting events exposes systematic inflation in the scheduled (median) forecast:
@@
(6) In safety-critical settings, OPR provides a simple operator-legible diagnostic for whether improvements in one-sided tail risk are achieved by trivial over-scheduling.
- Upward Reserve Requirement (Reserve).
@@
The additional upward capacity required to cover 99.5% of the under-forecast error distribution. We use the 99.5th percentile as a practical tail-risk proxy motivated by probabilistic adequacy practice (e.g., LOLE-style planning criteria; [11]). We report this both in megawatts (MW) and as a percentage of the point forecast (scale-free):
@@
(7) (8) Adding to the point forecast covers 99.5% of historical under-prediction events at the evaluated horizon. We report both MW and percentage: MW is normalization-free, while the percentage form is directly interpretable as an add-on to the scheduled point forecast. Because percentage-normalized reserve can be biased by systematic forecast inflation, we interpret Reserve jointly with UPR and explicit bias diagnostics (Bias24h / OPR) to ensure improvements do not come from trivial over-forecasting.
2.3.3 Motivating Example
@@
To illustrate the operational reality of these metrics, we examine empirical results (detailed later in Table 2) where tail-risk metrics and bias diagnostics move in different directions. For example, a model trained with a tail-focused loss might ostensibly reduce reserve requirements (Reserve), but do so by simply inflating the entire forecast—increasing Bias24h and Over-Prediction Rate (OPR) to unacceptable levels. This motivates our core reporting principle: any improvement in one-sided risk metrics must be interpreted jointly with bias diagnostics to avoid “fake safety” via systematic inflation.
2.3.4 Differentiable Surrogates for Risk Metrics
@@
While UPR and tail reserve requirements are excellent for evaluation, they are non-differentiable and difficult to optimize directly. To bridge this gap, we employ a Bias-Constrained Probabilistic Objective (detailed in Section 4.3) as a differentiable surrogate during training. This objective combines a weighted multi-quantile pinball loss to learn a calibrated distribution with explicit hinge penalties on bias and over-prediction rate to prevent systematic forecast inflation.
@@
In principle, emphasizing upper-tail quantiles increases the gradient penalty on under-predictions relative to over-predictions, aligning learning with asymmetric operational costs. However, as we demonstrate in Section 6, single-quantile training can be “gamed”; our constrained formulation (Section 4.3) ensures that improvements in tail risk do not come at the expense of acceptable schedule bias.
2.4 Value-Oriented Learning
@@
Traditional forecasting objectives (MSE/MAPE) operate under the assumption that prediction accuracy proxies decision quality. However, in grid operations, the cost function is highly asymmetric and constrained. The “Value-Oriented” or “Smart Predict-then-Optimize” paradigm [12] challenges the separation of prediction and downstream optimization, proposing that models should be trained to minimize the final operational cost. While theoretically optimal, full end-to-end bilevel optimization is often computationally intractable for large-scale deep learning. In this work, we adopt a scalable realization of this philosophy: instead of embedding the full extensive-form optimization solver, we encode the operational value function directly into a bias-constrained probabilistic objective (Section 4.3).
3 Related Work
3.1 Deep Learning for Time Series
Deep learning has systematically displaced statistical methods (ARIMA, SVR) for short-term load forecasting due to its superior ability to model non-linear relationships. Current research focuses on two primary paradigms for long-context multivariate modeling:
@@ 3.1 Deep Learning for Time Series
Channel Independence: Models like PatchTST [13] treat multivariate time series as independent channels and segment them into patches. This reduces complexity and represents the current state-of-the-art for Transformer accuracy, though it sacrifices explicit cross-variable correlation modeling.
@@
3.2 State Space Models for Time Series
@@
3.3 End-to-End and Value-Oriented Learning
@@
3.4 Weather-Integrated Forecasting
@@
3.5 Foundation Models for Time Series
Pre-trained foundation models have recently emerged as alternatives to task-specific architectures. Chronos [19] adapts language model pre-training to time series through quantization, demonstrating strong zero-shot performance on diverse forecasting benchmarks. However, the effectiveness of foundation models on domain-specific tasks with strong exogenous dependencies (e.g., weather-driven load) remains an open question, as these models typically lack mechanisms for incorporating auxiliary covariates.
4 Methodology
4.1 Model Architectures
We evaluate three Mamba-based architectures representing different design philosophies within the state space model paradigm, alongside the iTransformer baseline.
4.1.1 S-Mamba: The Linear-Time Baseline
@@
The primary goal of S-Mamba is to test whether the core selective state space mechanism alone—without complex decomposition or attention—is sufficient to model grid dynamics. Its design prioritizes computational efficiency ( complexity) above all else, making it an ideal candidate for resource-constrained edge deployment (0.08ms inference latency).
4.1.2 PowerMamba: Physics-Aware Decomposition
@@
Series Decomposition: It splits the input into "Trend" (low-frequency) and "Seasonal" (high-frequency) components, and processes them via independent Mamba encoders. This explicitly separates the "duck curve" drift from daily load cycles.
Bidirectionality: Unlike language modeling where causality is strict (future cannot affect past), time series analysis benefits from looking forward (e.g., smoothing). PowerMamba processes sequences in both forward and backward directions to capture a holistic temporal context.
4.1.3 Mamba-ProbTSF: Probabilistic Output Heads
Grid operators rarely make decisions based on a single deterministic number; they require confidence intervals to schedule reserves. Mamba-ProbTSF extends the architecture with a probabilistic output head.
@@ 4.1.3 Mamba-ProbTSF: Probabilistic Output Heads
This shift from minimizing MSE to maximizing likelihood enables the model to quantify "known unknowns," providing actionable risk metrics for uncertainty-aware dispatch.
@@
Alignment with our evaluations. In our main walk-forward results we evaluate this Gaussian-head variant as a probabilistic baseline; in the fixed-split loss-function ablations we fine-tune weather-initialized checkpoints with multi-quantile heads to study calibration and Bias/OPR-constrained objectives (Section 4.3, Table 2).
4.1.4 iTransformer: The Strongest Multivariate Baseline
The iTransformer [14] challenges the standard Transformer paradigm. Instead of tokenizing time steps (embedding steps into vectors), it tokenizes variables (embedding the entire time series of a variate as a single token).
Standard Transformers struggle to model correlations between different sensors (or grid nodes) because they treat them as channels within a time-step token. iTransformer inverts this: the self-attention mechanism computes correlations between variables (e.g., Load vs. Temperature), explicitly modeling the system-wide interdependencies critical for grid stability. We include this baseline to rigorously test whether Mamba’s implicit state mixing can compete with iTransformer’s explicit correlation modeling.
4.2 Weather-Integrated Architectures
@@
Motivated by the dependency of load on weather (analyzed in Section 6), we developed weather-integrated variants of each architecture. The key challenge is fusing meteorological covariates (temperature, humidity, solar radiation) with the load signal while respecting building thermal lag—HVAC systems respond to temperature changes with delays of 2–6 hours depending on building mass.
Fusion Strategies. S-Mamba uses Early Fusion (Fig. 1(a)): load and weather are concatenated at the input, requiring a wider embedding layer. PowerMamba uses Summation Fusion (Fig. 1(b)): weather features are added to both decomposed streams. Mamba-ProbTSF also uses summation fusion (Fig. 1(c)) with careful normalization to preserve uncertainty estimates. iTransformer uses Tokenization Fusion (Fig. 1(d)): each weather variable becomes a separate token, enabling cross-variable attention at the cost of increased complexity.
4.3 Bias-Constrained Probabilistic Objective

### v6

@@ Abstract
Using this framework, we evaluate five neural architectures—two state space models (S-Mamba, PowerMamba), two Transformers (iTransformer, PatchTST), an LSTM, and a probabilistic SSM variant (Mamba-ProbTSF)—on a weather-aligned California Independent System Operator (CAISO) dataset spanning Nov 2023–Nov 2025 (84,498 hourly records across 5 regional transmission areas) under a rolling-origin walk-forward backtest. We develop and evaluate thermal-lag-aligned weather fusion strategies matched to each architecture’s inductive bias.
Our results demonstrate that standard accuracy metrics are insufficient proxies for operational safety: models with comparable MAPE can imply materially different tail reserve requirements (Reserve). We show that explicit weather integration narrows error distributions, with the magnitude of improvement being architecturally determined—iTransformer’s cross-variate attention benefits significantly more than PatchTST’s channel-independent design. Crucially, we identify a widespread susceptibility to “fake safety” in risk-averse forecasting: while probabilistic calibration reduces upper-tail errors, it achieves this by systematically inflating schedules (e.g., increasing bias by over 1,700 MW in severe cases) if left unconstrained. To solve this, we introduce Bias/OPR-constrained objectives that enable auditable trade-offs between minimizing tail risk and preventing trivial over-forecasting.
I Introduction
@@
We present a systematic evaluation of state space model and Transformer architectures for reliable California grid load forecasting. Benchmarking Mamba variants against PatchTST, iTransformer, LSTM, and the foundation model Chronos, we shift the evaluation focus from pure accuracy to operational safety. Our walk-forward results highlight that accuracy alone is not a sufficient proxy for operational risk: Reserve can vary substantially across models even when MAPE is comparable. Furthermore, we demonstrate that explicit weather integration is critical for safety by narrowing error distributions and reducing extreme error events. Crucially, our loss-function ablations reveal a latent danger in risk-averse forecasting: unconstrained probabilistic training achieves apparent reductions in tail risk by systematically inflating schedules (e.g., driving positive bias from +109 MW to +1,862 MW). We demonstrate how to prevent this “fake safety” using explicit operational constraints.
@@
Grid-Specific Evaluation Framework. We formalize operational risk metrics (Under-Prediction Rate, tail Reserve requirements, Bias/OPR diagnostics) that capture asymmetric operational costs and prevent “fake safety” via systematic forecast inflation (Section II-C).
@@
Exposing “Fake Safety” via Constrained Objectives. We demonstrate that naive probabilistic calibration (multi-quantile pinball) can trivially reduce large-error events by introducing massive systematic inflation. We introduce and evaluate Bias/OPR constraints to enforce auditable trade-offs between true tail-risk reduction and schedule inflation (Section VI).
II Background
II-A State Space Models
@@
II-B Load Forecasting Fundamentals
@@
II-C Grid-Specific Evaluation Metrics
@@
II-C1 Operational Cost Asymmetry
@@
II-C2 Proposed Metrics
@@
Under-Prediction Rate (UPR). The frequency of under-estimation events represents the probability of needing real-time upward dispatch:
@@
A naive model might achieve low MAPE by constantly under-predicting (taking the median); UPR exposes this risky behavior.
@@
Over-Prediction Rate (OPR). The frequency of over-forecasting events exposes systematic inflation in the scheduled (median) forecast:
@@
In safety-critical settings, OPR provides a simple operator-legible diagnostic for whether improvements in one-sided tail risk are achieved by trivial over-scheduling.
@@
Upward Reserve Requirement (Reserve). The additional upward capacity required to cover 99.5% of the under-forecast error distribution. We use the 99.5th percentile as a practical tail-risk proxy motivated by probabilistic adequacy practice (e.g., LOLE-style planning criteria; [11]). We report this both in megawatts (MW) and as a percentage of the point forecast (scale-free):
@@
Adding to the point forecast covers 99.5% of historical under-prediction events at the evaluated horizon. We report both MW and percentage: MW is normalization-free, while the percentage form is directly interpretable as an add-on to the scheduled point forecast. Because percentage-normalized reserve can be biased by systematic forecast inflation, we interpret Reserve jointly with UPR and explicit bias diagnostics (Bias24h / OPR) to ensure improvements do not come from trivial over-forecasting.
II-C3 Motivating Example
@@
To illustrate the operational reality of these metrics, we examine empirical results (detailed later in Table II) where tail-risk metrics and bias diagnostics move in different directions. For example, a model trained with a tail-focused loss might ostensibly reduce reserve requirements (Reserve), but do so by simply inflating the entire forecast—increasing Bias24h and Over-Prediction Rate (OPR) to unacceptable levels. This motivates our core reporting principle: any improvement in one-sided risk metrics must be interpreted jointly with bias diagnostics to avoid “fake safety” via systematic inflation.
II-C4 Differentiable Surrogates for Risk Metrics
@@
While UPR and tail reserve requirements are excellent for evaluation, they are non-differentiable and difficult to optimize directly. To bridge this gap, we employ a Bias-Constrained Probabilistic Objective (detailed in Section IV-D) as a differentiable surrogate during training. This objective combines a weighted multi-quantile pinball loss to learn a calibrated distribution with explicit hinge penalties on bias and over-prediction rate to prevent systematic forecast inflation.
@@
In principle, emphasizing upper-tail quantiles increases the gradient penalty on under-predictions relative to over-predictions, aligning learning with asymmetric operational costs. However, as we demonstrate in Section VI, single-quantile training can be “gamed”; our constrained formulation (Section IV-D) ensures that improvements in tail risk do not come at the expense of acceptable schedule bias.
II-D Value-Oriented Learning
@@
Traditional forecasting objectives (MSE/MAPE) operate under the assumption that prediction accuracy proxies decision quality. However, in grid operations, the cost function is highly asymmetric and constrained. The “Value-Oriented” or “Smart Predict-then-Optimize” paradigm [12] challenges the separation of prediction and downstream optimization, proposing that models should be trained to minimize the final operational cost. While theoretically optimal, full end-to-end bilevel optimization is often computationally intractable for large-scale deep learning. In this work, we adopt a scalable realization of this philosophy: instead of embedding the full extensive-form optimization solver, we encode the operational value function directly into a bias-constrained probabilistic objective (Section IV-D).
III Related Work
III-A Deep Learning for Time Series
Deep learning has systematically displaced statistical methods (ARIMA, SVR) for short-term load forecasting due to its superior ability to model non-linear relationships [13, 14]. The power systems community has actively embraced deep neural networks to manage the increasing stochasticity of smart grids. Recent innovations include deep residual networks [15], densely connected networks [16], pooling-based deep RNNs [17], and ensemble learning frameworks [18, 19] for deterministic structural forecasting. To capture uncertainty, researchers have proposed Bayesian deep learning [20], probabilistic multitask prediction models [21], transfer-learning frameworks for masked behind-the-meter loads [22, 23, 24], collaborative multi-district approaches [25], and advanced quantile-mixing methods [26]. While these bespoke architectures are powerful, they typically rely on symmetric loss functions, motivating the need for our proposed Bias-Constrained Probabilistic Objective designed for asymmetric operational risk.
Current research on foundation architectures focuses on two primary paradigms for long-context multivariate modeling:
@@ III-A Deep Learning for Time Series
Channel Independence: Models like PatchTST [27] treat multivariate time series as independent channels and segment them into patches. This reduces complexity and achieves leading accuracy among Transformer architectures, though it sacrifices explicit cross-variable correlation modeling.
@@
III-B State Space Models for Time Series
@@
III-C End-to-End and Value-Oriented Learning
@@
III-D Weather-Integrated Forecasting
@@
III-E Foundation Models for Time Series
Pre-trained foundation models have recently emerged as alternatives to task-specific architectures. Models like Chronos [33] and TimesFM [34] adapt language model pre-training and decoder-only architectures to time series, demonstrating strong zero-shot performance on diverse forecasting benchmarks. However, the effectiveness of foundation models on domain-specific tasks with strong exogenous dependencies (e.g., weather-driven load) remains an open question, as these models typically lack mechanisms for incorporating auxiliary covariates.
IV Methodology
IV-A Model Architectures
We evaluate five neural architectures spanning three families—two state space models representing different SSM design philosophies, two Transformer variants with contrasting multivariate strategies, and an LSTM—alongside a probabilistic SSM variant (Mamba-ProbTSF) for uncertainty quantification.
IV-A1 S-Mamba: The Linear-Time Baseline
@@
S-Mamba tests whether the core selective state space mechanism alone—without complex decomposition or attention—suffices for grid forecasting. Its design prioritizes computational efficiency ( complexity), making it a candidate for resource-constrained edge deployment.
IV-A2 PowerMamba: Physics-Aware Decomposition
@@
Series Decomposition: It splits the input into “Trend” (low-frequency) and “Seasonal” (high-frequency) components, processing them via independent Mamba encoders. This explicitly separates the duck curve drift from daily load cycles.
Bidirectionality: PowerMamba processes sequences in both forward and backward directions to capture a holistic temporal context, since time series analysis (unlike causal language modeling) benefits from bidirectional information flow.
IV-A3 PatchTST: Channel-Independent Patching
PatchTST [27] represents the channel-independent paradigm for Transformer-based time series forecasting. Rather than tokenizing variables, PatchTST segments each univariate channel into fixed-length patches and processes them with a shared Transformer encoder. Channel independence means each variate shares the same Transformer weights, enabling efficient multi-channel processing without explicit cross-variable attention. This tests whether temporal pattern extraction alone suffices for accurate grid forecasting.
IV-A4 iTransformer: Cross-Variate Attention
The iTransformer [28] challenges the standard Transformer paradigm. Instead of tokenizing time steps (embedding steps into vectors), it tokenizes variables (embedding the entire time series of a variate as a single token). Self-attention therefore computes correlations between variables (e.g., Load vs. Temperature), explicitly modeling system-wide interdependencies. By design, this mechanism requires multiple variates; with a single load variate, self-attention reduces to near-identity.
IV-A5 Mamba-ProbTSF: Probabilistic Output Heads
Grid operators rarely make decisions based on a single deterministic number; they require confidence intervals to schedule reserves. Mamba-ProbTSF extends the S-Mamba architecture with a probabilistic output head.
@@ IV-A5 Mamba-ProbTSF: Probabilistic Output Heads
This shift from minimizing MSE to maximizing likelihood enables the model to quantify “known unknowns,” providing actionable risk metrics for uncertainty-aware dispatch.
@@
Alignment with our evaluations. In our main walk-forward results we evaluate this Gaussian-head variant as a probabilistic baseline; in the fixed-split loss-function ablations we fine-tune weather-initialized checkpoints with multi-quantile heads to study calibration and Bias/OPR-constrained objectives (Section IV-D, Table II).
IV-B Architecture Adaptations for Grid Forecasting
Direct application of generic time series models to load forecasting yields suboptimal results because these architectures lack two capabilities essential for grid operations: (1) awareness of temporal context—hourly and weekly demand cycles that dominate load profiles—and (2) the ability to selectively incorporate weather covariates whose influence on load is non-stationary and lagged.
We adapt each baseline with a shared set of domain-specific components while preserving the core architectural mechanism that each model is designed to test:
Temporal embeddings. All sequence models (SSMs, LSTM) receive learnable hour-of-day and day-of-week embeddings, concatenated and projected to , then added to the input representation at each time step. PatchTST processes temporal context implicitly through its patch structure; iTransformer encodes hour and day as additional variate tokens. These embeddings inject the strong 24h and 168h periodicities that characterize load.
Bidirectional encoding. SSMs and LSTM use bidirectional processing—separate forward and backward encoders whose outputs are fused via a learned linear projection. Bidirectional processing is applied strictly within the fixed-length lookback window ( hours); no information from the forecast horizon is accessible to the encoder.
Attention-weighted pooling. Rather than mean-pooling the encoder’s hidden states, SSMs and LSTM use a learned attention vector () to produce a weighted summary , allowing the model to selectively emphasize time steps most relevant to the forecast.
Modular weather integration. Each architecture receives weather covariates through a fusion mechanism matched to its inductive bias (detailed in Section IV-C). The weather integration layer is architecturally decoupled from the core encoder: it can be toggled on or off without retraining.
IV-C Weather-Integrated Architectures
@@
Motivated by the dependency of load on weather (analyzed in Section VI), we developed weather-integrated variants of each architecture. The key challenge is fusing meteorological covariates (temperature, humidity, solar radiation) with the load signal while respecting building thermal lag—HVAC systems respond to temperature changes with delays of 2–6 hours depending on building mass.
Fusion Strategies (Architecture-Matched)
S-Mamba (Early Summation, Fig. 1(a)). Weather embeddings are summed with load and temporal tokens prior to the BiMamba stack, projecting all exogenous signals into a unified dimensional space. The selective state space mechanism then uses this fused representation to modulate its continuous transition matrices.
PowerMamba (Pre-Decomposition Fusion, Fig. 1(b)). Weather features are fused before the sequence is decomposed. Because meteorological events drive both multi-day trend shifts and rapid sub-daily HVAC cycling, early fusion allows the decomposition module to route weather-induced variance into both the seasonal and trend branches.
PatchTST (Interleaved Cross-Attention, Fig. 1(c)). To preserve PatchTST’s core channel-independence inductive bias, we patch load and weather channels independently. A dedicated cross-attention sublayer allows load patches to query weather patches, permitting the load representation to attend to relevant meteorological context without entangling the self-attention dynamics.
iTransformer (Variate Tokenization, Fig. 1(d)). Each weather covariate is embedded as a distinct variate token. These weather tokens are concatenated with load and temporal tokens, exposing them to the global cross-variate attention mechanism.
LSTM (Early Concatenation). Weather features are concatenated with load features at each timestep, providing the recurrent gating mechanisms simultaneous access to demand and meteorological forcing.
IV-D Bias-Constrained Probabilistic Objective